OpenStudy (barrelracing):

please help suppose y varies directly with x. write a direct variation equation that relates x and y then graph the equation y=7.5 when x=1.5

OpenStudy (barrelracing):

@countryboy97 @cimfla

OpenStudy (barrelracing):

@Lady.Liv1776 @MD175

OpenStudy (barrelracing):

@bry856

OpenStudy (bry856):

y = kx 7.5=(1.5)*k k=5

OpenStudy (barrelracing):

so the direct variation equation is 7.5=(1.5)*k

jimthompson5910 (jim_thompson5910):

bry856 has shown that k = 5 after solving y = kx for k So y = kx turns into y = 5x when you replace k with 5

jimthompson5910 (jim_thompson5910):

y = 5x y = 5*1.5 ... replace x with 1.5 y = 7.5 so the equation checks out

OpenStudy (barrelracing):

thank you soo much I understand now.

jimthompson5910 (jim_thompson5910):

you're welcome